OpenStudy (winsteria):

SOMEONE PLEASE HELP ME IF YOU DO I'LL GIVE YOU A MEDAL AND FAN Dustin is mixing concrete. A formula for concrete calls for 3/8 gal of water. Dustin wants to make 7/9 more concrete than the formula creates. How much water should Dustin use? Express your answer as a fraction in simplest form.

OpenStudy (winsteria):

@TheSmartOne @mony01 @StudyGurl14

OpenStudy (studygurl14):

\(\large\frac{3}{8}+\frac{3}{8}(\frac{7}{9})\tt =the~answer\)

OpenStudy (winsteria):

oh okay thank you